0.24, calculate the following: If X follows Bernoulli distribution B, p > 0.5 and V(X) .48 .48 ELX] 48

The 2 guesses are incorrect. Please help to solve correct answer
Answer #1

Ans:

For Bernoulli distribution:

V(X)=p(1-p)=0.24

So,

p=0.6

1-p=1-0.6=0.4

E(X)=p=0.6